OpenStudy (anonymous):

Solve: 5(x + 7)2 = 25

OpenStudy (anonymous):

is that squared?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

5(x+7)*2 =25 / 5 --- (x+7)*2 =5 --- x*2 + 14x +49 =5 --- x*2 +14x +44 =0

jhonyy9 (jhonyy9):

14 +/- (196-88)*(1/2) 14 +/- (108)*(1/2) 14 +/- 2 radical 27 x_1,_2 = --------------------- = ------------------ = ------------------ 2 2 2 = 7 +/- radical27

jhonyy9 (jhonyy9):

= 7 +/- 3 radical 3

OpenStudy (anonymous):

5x^2 +70x +245=25 5x^2+70x+220 x=-7 - sqrt 5 x=sqrt 5 -7